10 points References 28
8 b. Find A5.5 sx 5)·(Round z value to 2 decimal places and final answer to 4 decimal places.) P(5.5 sXs7.5) c. Find x such that RX>x) 00594 (Round z value and final answer to 3 decimal places.) Print d. Find x such that Px s Xs3.4)-0.1255. (Negative value should be Indicated by a minus sign. Round 3 decimal places.) value and final answer to < Prev 8 of 12 Next > 28
Answer #1

a)

probability = P(X>6.5) = P(Z>1.55)= 1-P(Z<1.55)= 1-0.9394= 0.0606

b)

probability = P(5.5<X<7.5) = P(1.05<Z<2.05)= 0.9798-0.8531= 0.1267

c)

here z=1.56

hence x=3.4+1.56*2=6.520

d)

x=2.760